OpenStudy (srtauser):

How do you do this math problem: x/x^2-y^2 + y/x^2-y^2 answer options: A. 1/x-y B. 1/x+y C. -1/x+y D. 1/y-x

OpenStudy (zyberg):

Do you know \[\frac{ a }{ b } + \frac{ d }{ c } = \frac{ ac + db}{ bc }\]?

OpenStudy (srtauser):

no Thank you again! So it's B??

OpenStudy (welshfella):

No can you factor x^2 - y^2 ?

OpenStudy (welshfella):

the result of the addition is x + y --------- x^2 - y^2 If we factor the denominator we can simplify

OpenStudy (srtauser):

ok Its C

OpenStudy (srtauser):

I though when you simplify x^2-y^2 you get (x+y)(x-y)

OpenStudy (welshfella):

I think you are guessing x^2 - y^2 is an example of the difference of 2 squares x^2 - y^2 = ( x + y)( x - y)

OpenStudy (welshfella):